# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 2.12 Nonlocal models in pattern formations
In pattern formation, morphogenesis is a process by which spatial patterns and structures arise during the development of organisms gierer1972 . The creation of the various and complex forms found in organisms depends on this complex process. It involves a cascade of molecular and cellular processes that coordinate the orientation, polarity, and assembly of cells to generate distinct morphologies and configurations. Examples include (i) the development of an embryo into a fully formed organism, (ii) the differentiation process by which cells in developing organisms take on certain destinies and functions, and (iii) precise control of gene expression. The multidisciplinary study of morphogenesis aims to comprehend the fundamental processes that give rise to the astounding diversity of biological forms. It draws on the disciplines of genetics, developmental biology, and biophysics, among others.
@@
Spatiotemporal patterns in epidemiology reveal the dynamic interplay of place, time, and disease distribution within populations. Environmental, social, and biological factors all have an impact on these patterns tadic2023 ; mitrovic2023 . Epidemics can follow seasonal trends, exhibit cyclic behaviour, or demonstrate irregular patterns driven by a variety of factors, such as host susceptibility and environmental conditions, which are referred to as temporal patterns. However, spatio-temporal models deal with many spatial heterogeneity factors, such as higher or lower incidence rates in certain areas elliot2000 .
@@
Nonlocal models are essential for describing developing patterns in a variety of biological systems, especially when it comes to cell communication, cell adhesion, and bacterial aggregation. Cell adhesion models take into account distant adhesive forces, enabling cells to recognize and respond to the spatial distribution of adhesion molecules armstrong2006 . It can capture the collective behaviour of bacteria by addressing long-range interactions, such as quorum sensing, where bacterial cells communicate and coordinate their activities across longer distances miller2001 .
@@
Nonlocal models derived from quantum confinement are intriguing in characterizing pattern generation at the nanoscale and have several applications in materials research, quantum mechanics, and device manufacturing. Quantum confinement is the phenomenon in which the electronic and optical characteristics of materials are significantly affected by their size, particularly when limited to dimensions equivalent to electrons’ characteristic length scale otten2003 . The nonlocal models derived from quantum mechanics are frequently used to characterize the electrical and optical behaviour of quantum dots, nanowires, and other nanostructures michalek2015 ; taghipour2017 . The discrete energy levels and wave functions of confined electrons produce distinct patterns in their characteristics. Quantum dot solar cells, light-emitting diodes, and quantum dot transistors are some of the uses schornbaum2015 ; kahmann2020 . Nonlocal models can assist in explaining the development of energy bands and electronic patterns in quantum dot arrays, which influence their optical and electrical characteristics freeney2022 . Furthermore, quantum confinement can cause nonlocal electron interactions and transport behaviours, affecting the operation of nanoelectronic devices such as sensors, catalysis, and quantum computing.
@@
An interesting perspective on the emergence of complex patterns in many physical and biological systems may be gained from nonlocal models that describe pattern development resulting from symmetry-breaking processes ishihara2018 ; denk2020 . Symmetry-breaking is a basic idea in which complex patterns are formed when a system changes from a symmetrical to an asymmetrical state. These occurrences, prompted by disturbances or outside factors, start to establish patterns. Fluid dynamics, chemical reactions, and biological processes are examples of systems where this may happen crawford1991 ; islas2005 ; li2010 ; tanzy2013 ; mcglinchey2022 .
@@ 2.13 Complex nonlocal systems, multi-stability, and control
Nonlocal interactions enhance the species’ coexistence in the Lotka-Volterra competition model, and it supports how two competing species can coexist in the same resource hutchinson1961 . Different mechanisms have been proposed to explain such coexistence in ecological communities with the competitional environment levins1971 ; chesson2000 ; amarasekare2003 ; kartzinel2015 ; maciel2018 . Authors in maciel2021 proposed a nonlocal competition model as
@@
where $\Phi_{ij}$ is a kernel function and represents the influence of $j$-individuals on the $i$-individual when they are at a distance $|x-y|$ jdmurray . In this case, a normalized top-hat kernel function can be considered, and the authors in maciel2021 presented differently to incorporate spatial scales. For two species, it is given by:
$\Phi_{ij}(r)=\Pi_{2R_{ij}}=\begin{cases}\frac{1}{2R_{ij}}\leavevmode\nobreak\ \leavevmode\nobreak\ r\leq R_{ij}\\ \leavevmode\nobreak\ \leavevmode\nobreak\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{otherwise},\end{cases}$
where $R_{ij}$ is the interaction range and is a function of the species-specific ranges of competition, $R_{i}$ and $R_{j}$. Depending on these spatial scales $R_{i}$ and $R_{j}$, two possibilities can occur: additive and non-additive influence ranges. For the additive influence ranges, both the ranges of the competition overlap with each other tarnita2017 . In this case, the range of interspecific competition is $R_{i}+R_{j}$, and the ranges of intra-specific competitions are $2R_{i}$ and $2R_{j}$. This leads to a change in the kernel function ($\Phi_{ij}(r)=\Phi_{ji}(r)=\Pi_{2(R_{i}+R_{j})}(r)$, $\Phi_{ii}(r)$ = $\Pi_{2R_{i}}(r)$, and $\Phi_{jj}(r)$ = $\Pi_{2R_{j}}(r)$). On the other hand, for non-overlapping ranges of competitions bais2003 ; granato2019 , the kernel function remains the same for inter- and intra-specific competitions ($\Phi_{ij}(r)=\Pi_{2R_{i}}(r)$, $\Phi_{ji}(r)=\Pi_{2R_{j}}(r)$, $\Phi_{ii}(r)$ = $\Pi_{2R_{i}}(r)$, and $\Phi_{jj}(r)$ = $\Pi_{2R_{j}}(r)$).
A mixed hyperbolic-parabolic problem is studied in the prey-predator model in accounting for the flow of predators’ direction colombo2014 . We consider the mathematical problem as
@@
$\frac{\partial v}{\partial t}-\nabla\cdot(v\xi(u))=(\alpha u-\beta)v,$
where $u(\mathbf{x},t)$ and $v(\mathbf{x},t)$ are the prey and predator densities, respectively, at the position $\mathbf{x}\in\Omega\subset\mathbb{R}^{n}$ and time $t\in\mathbb{R}^{+}$. Here, the term $v\xi(u)$ represents the flow preferred to predators’ direction, and the velocity function $\xi(u)$ is generally nonlocal and nonlinear. A typical choice is colombo2014 :
@@
Here, the denominator $\sqrt{1+||\nabla(\phi\ast u)||^{2}}$ is the normalization factor. Prey species spread in all directions in this model, but predators have a direct movement that migrates towards higher prey density zones. Furthermore, the radius of support of $phi$ specifies how far predators can detect the presence of prey and, as a result, the direction in which they move.
@@
Bio-cells participate in complex, dynamic interactions encompassing various molecular, cellular, and tissue activities. These interactions are essential for living things to operate normally and have significance in several physiological and pathological situations. Cells, for example, communicate using complex signalling networks. Hormones, growth factors, and neurotransmitters connect to cell surface receptors, triggering a chain reaction of events within the cell alberts2017 . Several approaches are used to simulate and model these cellular processes, including metabolism. In flux balance analysis, we operate the cellular processes under steady-state conditions and use constraints based on mass balances, thermodynamics, and other factors to predict metabolic flux distributions orth2010 . The kinetic model involves individual reactions with kinetic reactions represented by ordinary or partial differential equations and studying how the concentrations change over time. In the cellular automata model, each cell follows some rules governing its interactions, and these rules determine its state, reflecting biological behaviours. In addition, agent-based models simulate individual agents (cells or molecules) and their interactions and help in studying emergent properties and spatial considerations.
@@
Nonlocal PDEs with nonlinear diffusion terms are effective mathematical tools for understanding nonlocal interactions and complex diffusion processes. These equations extend typical diffusion models by taking into account impacts from distant locations. Nonlocal PDEs with nonlinear diffusion are used to represent developed patterns in biological systems, such as population dispersion or the generation of spatial patterns in ecosystems. Complexity and pattern production in biophysical systems is a diverse field that spans several scales, from molecular interactions to tissue and organism-level structures wolpert2015 . Furthermore, microbial systems are exceptionally diversified and complex, ranging from intricate interactions within microbial communities to involvement in environmental processes and human health madsen2015 .
@@
The modelling of many-particle systems is also a complex task that has applications in physics, chemistry, biology, and a variety of other sciences. Newton’s equations of motion are frequently used in classical physics to describe the motion of particles in a many-body system goldstein2002 . In addition, statistical mechanics gives a statistical description of a vast number of particles feynman2018 . In kinetic theory, mathematical models involve the construction of equations that explain particle distribution and behaviour. Furthermore, network and graph-based models give a solid foundation for understanding the structure, dynamics, and behaviour of complex interactions in a variety of systems melnik2012 ; antoniouk2013 ; melnik2015 .
@@
Self-organized criticality (SOC) has been studied in a variety of disciplines, including evolutionary biology and dynamic models with complex relationships jensen1998 ; tadic2017 ; tadic2021a ; tadic2023 . SOC refers to the tendency of complex systems to grow towards a critical state in which minor events can lead to large-scale cascades and emergent behaviour, e.g., in the forest-fire model grassberger1993 , in the brain plenz2021 , etc. Complex systems naturally evolve to a critical state, and SOC provides a valuable framework for understanding the dynamics of evolutionary and complex interactive processes.
@@
Complex systems comprise a large number of linked components with nonlinear interactions, which frequently results in emergent phenomena that cannot be easily derived from individual components. When just a subset of the system’s variables or states are available for observation, it introduces uncertainty in precisely understanding and forecasting the system’s behaviour rupe2022 . The integration of these concepts is critical for acquiring insights into real-life occurrences that are characterized by both partial observations and complex interactions. Modelling such systems is an interdisciplinary process that requires an understanding of mathematics, statistics, and the specific application domain melnik1997 ; melnik1998a .
@@
Control problems involving human variables are common in many fields, including automation, transportation, and healthcare, where humans interact with complex systems. These systems incorporate human factors that examine cognitive, perceptual, and physical elements of human interaction with technology russell2019 . The objective is to create systems that take into account human skills, limits, and behaviours, maximizing performance and reducing errors. Additionally, these models are used in neurorehabilitation and robotics to combine muscle mechanics and control, which leads logically to planning and behaviour burdet2013 .
@@
Nonlocal high-order connectivity and interactions are the studies of long-distance effects and correlations between distant locations in a system, with an emphasis on higher-order links beyond close neighbours. This notion is especially important in a wide range of scientific areas, from physics and materials science to neurology and complex systems analysis battiston2020 ; bick2023 . Exploration of nonlocal high-order connectivity in the context of neurodegenerative diseases has become increasingly important in understanding the complex dynamics of brain function, cognition, information processing, and the progression of conditions such as Alzheimer’s, Parkinson’s, and others pal2022sr ; herzog2022 .

### v2

@@ 2.11 Nonlocal models in pattern formations
In reaction-diffusion systems such as the Gierer-Meinhardt system, a limiting situation known as the “Shadow System” occurs when one component (usually the inhibitor) diffuses significantly faster than the other [179, 180, 181]. In this case, the fast-diffusing component instantly enters a quasi-steady state, resulting in a reduced system closely matching the original system’s behaviour. Researchers have shown that the Gierer–Meinhardt system can be controlled by only the dynamics of the activator itself given by the nonlocal problem [182, 183]:
[…]
with positive initial and no-flux boundary conditions. Furthermore, the authors obtained global-in-time existence and blow-up solutions for the nonlocal problem in both finite and infinite time. The nonlocal model predicts the regulated expansion of the activator, which is ensured by global-in-time existence results; yet, the activator’s unlimited growth is relevant to finite- and infinite-time blow-up results.
@@
In pattern formation, morphogenesis is a process by which spatial patterns and structures arise during the development of organisms [184]. The creation of the various and complex forms found in organisms depends on this complex process. It involves a cascade of molecular and cellular processes that coordinate their orientation, polarity, and assembly of cells to generate distinct morphologies and configurations. Examples include (i) the development of an embryo into a fully formed organism, (ii) the differentiation process by which cells in developing organisms take on certain destinies and functions, and (iii) precise control of gene expression. The multidisciplinary study of morphogenesis aims to understand the fundamental processes that give rise to the astounding diversity of biological forms. It draws on the disciplines of genetics, developmental biology, and biophysics, among others.
@@
Spatio-temporal patterns in epidemiology reveal the dynamic interplay of place, time, and disease distribution within populations. Environmental, social, and biological factors all have an impact on these patterns [185, 186]. Epidemics can follow seasonal trends, exhibit cyclic behaviour, or demonstrate irregular patterns driven by various factors, such as host susceptibility and environmental conditions, referred to as temporal patterns. However, spatio-temporal models deal with many spatial heterogeneity factors, such as higher or lower incidence rates in certain areas [187].
@@
Nonlocal models are essential for describing the development patterns in various biological systems, especially when it comes to cell communication, cell adhesion, and bacterial aggregation. Cell adhesion models take into account distant adhesive forces, enabling cells to recognize and respond to the spatial distribution of adhesion molecules [154, 188]. It can capture the collective behaviour of bacteria by addressing long-range interactions, such as quorum detection, where bacterial cells communicate and coordinate their activities over longer distances [189].
@@
Nonlocal models derived from quantum confinement are intriguing in characterizing pattern generation at the nanoscale and have several applications in materials research, quantum mechanics, and device manufacturing. Quantum confinement is the phenomenon in which the electronic and optical characteristics of materials are significantly affected by their size, particularly when limited to dimensions equivalent to the characteristic length scale of electrons [190]. The nonlocal models derived from quantum mechanics are frequently used to characterize the electrical and optical behaviour of quantum dots, nanowires, and other nanostructures [191, 192]. The discrete energy levels and wave functions of the confined electrons produce distinct patterns in their characteristics. Quantum dot solar cells, light-emitting diodes, and quantum dot transistors are some of the uses [193, 194]. Nonlocal models can assist in explaining the development of energy bands and electronic patterns in quantum dot arrays, which influence their optical and electrical characteristics [195]. Furthermore, quantum confinement can cause nonlocal electron interactions and transport behaviours, affecting the operation of nanoelectronic devices such as sensors, catalysis, and quantum computing.
@@
An interesting perspective on the emergence of complex patterns in many physical and biological systems may be gained from nonlocal models that describe pattern development resulting from symmetry-breaking processes [196, 197]. Symmetry-breaking is a basic idea in which complex patterns are formed when a system changes from a symmetrical to an asymmetrical state. These occurrences, prompted by disturbances or outside factors, start to establish patterns. Fluid dynamics, chemical reactions, and biological processes are examples of systems where this may happen [198, 199, 200, 201, 202].
@@ 2.12 Complex nonlocal systems, multi-stability, and control
Nonlocal interactions enhance the species’ coexistence in the Lotka-Volterra competition model, and it supports how two competing species can coexist in the same resource [203]. Different mechanisms have been proposed to explain such coexistence in ecological communities with the competitional environment [204, 205, 206, 207, 208]. Authors in [209] proposed a nonlocal competition model as
@@
where $\Phi_{ij}$ is a kernel function and represents the influence of $j$-individuals on $i$-individuals when they are at a distance $|x-y|$ [4]. In this case, a normalized top-hat kernel function can be considered, and the authors in [209] presented differently to incorporate spatial scales. For two species, it is given by:
$\Phi_{ij}(r)=\Pi_{2R_{ij}}=\begin{cases}\frac{1}{2R_{ij}}&r\leq R_{ij}\\ \leavevmode\nobreak\ \leavevmode\nobreak\ 0&\mbox{otherwise},\end{cases}$
where $R_{ij}$ is the interaction range and is a function of the species-specific ranges of competition, $R_{i}$ and $R_{j}$. Depending on these spatial scales $R_{i}$ and $R_{j}$, two possibilities can occur: additive and non-additive influence ranges. For the additive influence ranges, both the ranges of the competition overlap with each other [210]. In this case, the range of interspecific competition is $R_{i}+R_{j}$, and the ranges of intra-specific competitions are $2R_{i}$ and $2R_{j}$. This leads to a change in the kernel function ($\Phi_{ij}(r)=\Phi_{ji}(r)=\Pi_{2(R_{i}+R_{j})}(r)$, $\Phi_{ii}(r)$ = $\Pi_{2R_{i}}(r)$, and $\Phi_{jj}(r)$ = $\Pi_{2R_{j}}(r)$). On the other hand, for non-overlapping ranges of competitions [211, 212], the kernel function remains the same for inter- and intra-specific competitions ($\Phi_{ij}(r)=\Pi_{2R_{i}}(r)$, $\Phi_{ji}(r)=\Pi_{2R_{j}}(r)$, $\Phi_{ii}(r)$ = $\Pi_{2R_{i}}(r)$, and $\Phi_{jj}(r)$ = $\Pi_{2R_{j}}(r)$). The range of the kernel function is crucial to the nonlocal competition. A narrow range shows that competition is mostly local, while a wide range suggests that individuals compete over a broader area. These types of nonlocal models produce stable spatial structures similar to physical systems where repulsive forces lead to the separation of particles. In addition, they are used to study phenomena such as species segregation, biodiversity, and the evolution of social behaviours.
A mixed hyperbolic-parabolic problem is studied in the predator-prey model in accounting for the flow of predators’ direction [213]. We consider the mathematical problem as
@@
$\frac{\partial v}{\partial t}+\nabla\cdot(v\xi(u))=(\alpha u-\beta)v,$
where $u(\mathbf{x},t)$ and $v(\mathbf{x},t)$ are the prey and predator densities, respectively, at the position $\mathbf{x}\in\Omega\subset\mathbb{R}^{n}$ and time $t\in\mathbb{R}^{+}$. Here, the term $v\xi(u)$ represents the flow preferred predators’ direction, and the velocity function $\xi(u)$ is generally nonlocal and nonlinear. A typical choice is [213]:
@@
Here, the denominator $\sqrt{1+||\nabla(\phi\ast u)||^{2}}$ is the normalization factor. Prey species spread in all directions in this model, but predators have a direct movement that migrates towards higher prey density zones. Furthermore, the radius of support of $\phi$ specifies how far predators can detect the presence of prey and, as a result, the direction in which they move.
Another type of parabolic nonlocal model where the nonlocal term appears as the master equation’s mean field limit when reaction radius is used, which is given by [214, 215, 216, 179]:
[…]
with non-negative initial and no-flux boundary conditions in a bounded domain $\Omega\subset\mathbb{R}^{n}$. Here, $q_{A}$ and $q_{B}$ are the probabilities of the molecules $A$ and $B$ located at position $\mathbf{x}$ and at time $t$, respectively, and $B(\mathbf{x},R)=\{\mathbf{y}\in\mathbb{R}^{n}:|\mathbf{y}-\mathbf{x}|<R\}$. In a chemical reaction, this $R$ represents the reaction radius of the fundamental reaction process. In this case, the assumption is that the chemical reaction occurs if and only if a pair of two different chemical particles is in the distance $R$. The mean-field limit approach is commonly used in statistical mechanics and kinetic theory, where it simplifies complex many-body systems by averaging interactions over a large number of particles [217].
In evolutionary game dynamics, a key component is to describe how a population of players adjusts their strategies throughout the game in response to the strategies’ level of success. Nonlocal models have gained prominence in studying such dynamics, notably replicator dynamics. These models account for interactions that are not only local neighbours but can also involve distant people or entire communities, which is critical for accurately portraying the complexity of real-world settings. For a finite-dimensional strategy space $S=\{1,2,\ldots,m\}$, the replicator dynamics equation can be written as [218, 219]:
${}\frac{dp_{i}(t)}{dt}=\bigg{(}\sum_{j=1}^{m}a_{ij}p_{j}(t)+\sum_{k=1}^{m}\sum_{j=1}^{m}a_{kj}p_{j}(t)p_{k}(t)\bigg{)}p_{i}(t),\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,\ldots,m,$ (27)
where $a_{ij}$ is the payoff for players 1 and 2 with strategies $i$ and $j$. Here, the set $p=(p_{1},\ldots,p_{m})$ satisfies the conditions $p_{i}\geq 0$, $i=1,\dots,m$, and $\sum_{i=1}^{m}p_{i}=1$. If the strategy space consists of an infinite number of strategies, then the replicator dynamics take the form:
${}\frac{dp_{i}(t)}{dt}=\bigg{(}\sum_{j\in\mathbb{Z}}a_{ij}p_{j}(t)+\sum_{k\in\mathbb{Z}}\sum_{j\in\mathbb{Z}}a_{kj}p_{j}(t)p_{k}(t)\bigg{)}p_{i}(t),\leavevmode\nobreak\ \leavevmode\nobreak\ i=1,2,\ldots,$ (28)
with $p_{i}(t)\geq 0$ and $\left\lVert p\right\rVert_{l^{1}}=1$. Furthermore, we can rewrite the replicator dynamics equation in terms of absolutely continuous Lebesgue measure with density $u(t,x)$ as:
[…]
where $f(x,y)$ is the payoff for players 1 and 2 play $x$ and $y$, respectively. Additional details can be found in [218].
@@
Bio-cells participate in complex, dynamic interactions encompassing various molecular, cellular, and tissue activities. These interactions are essential for living things to operate normally and have significance in several physiological and pathological situations. Cells, for example, communicate using complex signalling networks. Hormones, growth factors, and neurotransmitters connect to cell surface receptors, triggering a chain reaction of events within the cell [220]. Several approaches are used to simulate and model these cellular processes, including metabolism. In flux balance analysis, we operate the cellular processes under steady-state conditions and use constraints based on mass balances, thermodynamics, and other factors to predict metabolic flux distributions [221]. The kinetic model involves individual reactions with kinetic reactions represented by ordinary or partial differential equations and studying how the concentrations change over time. In the cellular automata model, each cell follows some rules governing its interactions, and these rules determine its state, reflecting biological behaviours. In addition, agent-based models simulate individual agents (cells or molecules) and their interactions and help in studying emergent properties and spatial considerations.
@@
Nonlocal PDEs with nonlinear diffusion terms are effective mathematical tools for understanding nonlocal interactions and complex diffusion processes. These equations extend typical diffusion models by accounting for impacts from distant locations, and they generate patterns in biological systems, such as population dispersion or the generation of spatial patterns in ecosystems. Complexity and pattern production in biophysical systems is a diverse field that spans several scales, from molecular interactions to tissue and organism-level structures [222]. Furthermore, microbial systems are exceptionally diversified and complex, ranging from intricate interactions within microbial communities to involvement in environmental processes and human health [223].
@@
Modelling many-particle systems is also complex and has applications in physics, chemistry, biology, and other sciences. Newton’s equations of motion are frequently used in classical physics to describe the motion of particles in a many-body system [224]. In addition, statistical mechanics gives a statistical description of a vast number of particles [225]. In kinetic theory, mathematical models involve the construction of equations that explain particle distribution and behaviour. Furthermore, network and graph-based models give a solid foundation for understanding the structure, dynamics, and behaviour of complex interactions in various systems [226, 227, 228].
@@
Self-organized criticality (SOC) has been studied in various disciplines, including evolutionary biology and dynamic models with complex relationships [229, 230, 231, 185]. SOC refers to the tendency of complex systems to grow towards a critical state in which minor events can lead to large-scale cascades and emergent behaviour, e.g., in the forest-fire model [232], in the brain [233], etc. Complex systems naturally evolve to a critical state, and SOC provides a valuable framework for understanding the dynamics of evolutionary and complex interactive processes.
@@
Complex systems comprise many linked components with nonlinear interactions, frequently resulting in emergent phenomena that cannot be easily derived from individual components. When just a subset of the system’s variables or states are available for observation, it introduces uncertainty in precisely understanding and forecasting the system’s behaviour [234]. The integration of these concepts is critical for acquiring insights into real-life occurrences that are characterized by both partial observations and complex interactions. Modelling such systems is an interdisciplinary process that requires understanding mathematics, statistics, and the specific application domain [235, 236].
@@
Control problems involving human variables are common in many fields, including automation, transportation, and healthcare, where humans interact with complex systems. These systems incorporate human factors that examine cognitive, perceptual, and physical elements of human interaction with technology [237]. The objective is to create systems that account for human skills, limits, and behaviours, maximizing performance and reducing errors. These models are also used in neurorehabilitation and robotics to combine muscle mechanics and control, which logically leads to planning and behaviour [238].
@@
Nonlocal high-order connectivity and interactions are the studies of long-distance effects and correlations between distant locations in a system, emphasising higher-order links beyond close neighbours. This notion is especially important in a wide range of scientific areas, from physics and materials science to neurology and complex systems analysis [239, 240]. Exploration of nonlocal high-order connectivity in the context of neurodegenerative diseases has become increasingly important in understanding the complex dynamics of brain function, cognition, information processing, and the progression of conditions such as Alzheimer’s, Parkinson’s, and others [46, 241, 242].